### v7

# Turning Up the Heat: Min- Sampling for Creative and Coherent LLM Outputs

Abstract
Large Language Models (LLMs) generate text by sampling the next token from a probability distribution over the vocabulary at each decoding step. Popular sampling methods like top- (nucleus sampling) often struggle to balance quality and diversity, especially at higher temperatures which lead to incoherent or repetitive outputs. We propose min- sampling, a dynamic truncation method that adjusts the sampling threshold based on the modelβs confidence by using the top tokenβs probability as a scaling factor. Our experiments on benchmarks including GPQA, GSM8K, and AlpacaEval Creative Writing show that min- sampling improves both the quality and diversity of generated text across different model families (Mistral and Llama 3) and model sizes (1B to 123B parameters), especially at higher temperatures. Human evaluations show a preference for min-, in both text quality and creativity. Min- sampling has been adopted by popular open-source LLM frameworks, including Hugging Face Transformers, VLLM, and many others, highlighting its considerable impact on improving text generation quality.
1 Introduction
A central challenge for Large Language Models (LLMs) is the trade-off between creativity and coherence during text generation. Popular sampling strategies like top- sampling (Holtzman etΒ al., 2020) and temperature scaling (Ackley etΒ al., 1985) are widely adopted, but increasing temperature to enhance diversity often reduces coherence, while more conservative sampling limits creativity.
We address the quality-diversity tradeoff with min- sampling, a dynamic truncation method that adjusts the sampling threshold based on model confidence using the top tokenβs probability as a scaling factor. This approach includes diverse options when the model is uncertain and keeps only high-confidence tokens when confident, balancing creativity and coherence even at high temperature.
Our experiments on benchmarks including GPQA (Rein etΒ al., 2023), GSM8K (Cobbe etΒ al., 2021), and AlpacaEval Creative Writing (Li etΒ al., 2023b). Our results show that min- sampling outperforms top- and other popular decoding methods such as top- (Fan etΒ al., 2018) and -sampling (Hewitt etΒ al., 2022a) by maintaining coherence while allowing for increased diversity, particularly at high-temperatures. We also conducted a comprehensive human evaluation to compare the quality and diversity of text generated by min- with that generated by traditional sampling methods. The results indicate a clear preference for min-, with participants rating min- outputs as superior in quality and diversity.
The contributions of this paper are as follows:
β’
We introduce min- sampling, a novel dynamic truncation method that effectively balances creativity and coherence in LLM-generated text, particularly at high temperatures.
β’
We present comprehensive experimental results across benchmarks, model families and sizes, demonstrating that in practice min- consistently improves quality and diversity of generated text, especially at higher temperatures, compared to sampling methods like top-.
β’
We validate the practical utility of min- through an extensive human evaluation, showing instances where human evaluators prefer min- outputs over those generated by other methods in terms of both quality and diversity.
β’
We provide practical empirical guidelines for hyperparameter selection and usecases.
β’
The rapid adoption of min- by the open-source LLM community highlights its practical benefits, even for large reasoning models such as Deepseek R1 (DeepSeek-AI etΒ al., 2025).
By introducing min- and offering empirical guidelines for its use, we enable exploration of high-temperature creative text generation without sacrificing coherence. Our results demonstrate that min- is a viable and superior alternative to existing sampling methods, both at standard and high-temperature settings, respectively, marking a major contribution to generative language modeling.
2 Related Work
We review current sampling methods and quality-diversity tradeoff, establishing motivation for min-.
Greedy Decoding and Beam Search.
These deterministic decoding strategies select the highest probability token at each step (Freitag & Al-Onaizan, 2017), which often produces repetitive and generic text due to lack of diversity. Beam search also incurs a notable runtime performance penalty.
Stochastic Sampling Methods.
These methods inject diversity via randomness in token selection. Temperature scaling adjusts distribution sharpness to balance diversity and coherence (Ackley etΒ al., 1985); however, higher temperatures often give incoherent outputs, limiting applicability. Top- sampling selects from the most probable tokens (Fan etΒ al., 2018), offering a simple way to prevent sampling unlikely tokens, but doesnβt dynamically adapt to varying confidence levels across contexts.
Top- sampling (nucleus sampling) restricts the token pool to those whose cumulative probability exceeds a total threshold (Holtzman etΒ al., 2020), balancing quality and diversity by preserving the "nucleus" of high-probability tokens. However, at higher temperatures, top- can still allow low-probability tokens into the sampling pool, leading to incoherent outputs. This specific trade-off between creativity and coherence at high temperatures is a key limitation we address with min-.
Entropy-Based Sampling.
Recent approaches like (-sampling) and mirostat sampling dynamically adjust the sampling pool based on token distribution entropy (Hewitt etΒ al., 2022a; Basu etΒ al., 2021). While these show promise, they often require complex parameter tuning and can be computationally expensive. We detail our experimental challenges running sampling in SectionΒ A.3.
Dynamic Threshold Methods.
Recent work has explored adaptive threshold mechanisms that adjust based on model confidence. Li etΒ al. (2023a) introduced an βadaptive plausibility constraintβ in their contrastive decoding framework, which filters tokens using a threshold proportional to the maximum probability: . This constraint is similar to min- as a threshold relative to model confidence rather than absolute. The key distinction is that Li etΒ al. (2023a) developed this as a constraint within a contrastive objective using expert and amateur models, min- treats confidence-scaled truncation as the standalone sampling mechanism, a much simpler approach.
3 Min- Sampling
The core idea of min- sampling is to dynamically adjust the sampling threshold based on the modelβs confidence at each decoding step, improving sensitivity to context and uncertainty while better balancing creativity and coherence, especially at high temperatures.
3.1 Overview of Min- Sampling
In autoregressive generation, a language model predicts the probability distribution over the vocabulary for the next token, conditioned on the existing sequence. At each step, the model selects a token from this distribution. Min- sampling is a stochastic method that varies its truncation threshold based on the modelβs confidence, making the threshold context-sensitive.
Formally, at each time step , let denote the vocabulary, and represent the conditional probability distribution over the vocabulary for the next token . Min- sampling works as follows:
1.
Calculate the Maximum Probability: Identify the maximum probability token in the distribution, denoted as .
2.
Truncation Threshold: Set a base probability threshold, Scale it by :
(1) This threshold ensures that tokens with sufficiently high relative probabilities are considered while filtering out less probable tokens in a context-dependent manner.
3.
Define Sampling Pool: Construct pool of tokens with probabilities :
(2) -
4.
Sample from Pool: Sample the next token from the normalized set :
3.2 Intuition Behind Min- Sampling
The key intuition behind min- sampling is that token truncation thresholds are relative and depend on how certain the distribution is for that token, and not absolute thresholds. When the model is highly confident about the next token (i.e., is high), min- restricts the pool to high-probability candidates to maintain coherence. Conversely, when the model is less confident, relaxing the sampling pool allows for a more creative and diverse generation. Unlike top- sampling, which truncates the distribution based on a fixed cumulative probability, min- dynamically adjusts the threshold based on the modelβs confidence, leading to more context-sensitive generation.
Figure 1 illustrates the effects of different sampling methods, including min-, on token probability distributions. In subfigure (a), we show an initial probability distribution over tokens. Subfigures (b), (c), and (d) demonstrate how top-, top-, and min- sampling truncate this distribution. Min- dynamically adjusts its filtering threshold based on the modelβs confidence, focusing on high-probability tokens when confident and including more diverse options when uncertain. This dynamic behavior helps min- balance coherence and diversity more effectively than top- and top- sampling.
3.3 Advantages Over Existing Methods
Min- sampling provides several advantages over existing sampling methods:
Balancing Creativity and Coherence.
Unlike more static thresholds like top- and top- that may allow either overly diverse (incoherent) or conservative (repetitive) token choices in the token sampling pool, min-βs can dynamically adjust its sampling threshold to different contexts within the same generated sequence based on the modelβs confidence in its outputs.
Robustness at High Temperatures.
A key limitation of existing sampling methods is their performance at high temperatures. As the temperature increases, token probabilities become more uniform, allowing unlikely tokens to be selected, resulting in incoherent text. Min- addresses this by scaling the truncation threshold proportional to the modelβs confidence, preserving output coherence even at higher temperatures. This makes it particularly valuable for tasks that benefit from high creativity, such as storytelling or generating diverse solutions (Tajwar etΒ al., 2025).
Simplicity.
Min- is simple, requiring minimal calculations. Unlike methods involving auxiliary models (Li etΒ al., 2023a) or complex entropy-based adjustments (Hewitt etΒ al., 2022a), min- integrates easily into existing LLM inference pipelines without substantial overhead, offering practical advantages for both research and real-world applications compared to entropy-based methods such as and sampling (Hewitt etΒ al., 2022b), as we discuss in SectionΒ A.3.
3.4 Implementation Details
Min- requires minimal changes to standard LLM decoding pipelines. Here, we provide reference:
Integration into Decoding Pipelines.
Min- sampling has been implemented as a logits processor in frameworks like Hugging Face Transformers (Wolf etΒ al., 2020) and , vLLM (Kwon etΒ al., 2023). After applying temperature scaling, the scaled threshold is computed, and tokens with probabilities below this threshold are filtered out before sampling. These operations are efficiently implemented using vectorized computations, adding negligible overhead to the decoding process.
Parameter Selection Guidelines.
β’
Choosing the Base Threshold (): Setting between 0.05 to 0.1 effectively balances creativity and coherence across tasks and models. Higher values of (e.g., 0.5-0.7 or even close to 1) can help maintain coherence and improve benchmark scores at very high temperatures, typically at the cost of diversity (see TableΒ 13).
β’
Temperature Settings: Min- works well across a wide range of temperatures, with higher temperatures (e.g., or ) enhancing diversity with much lower loss of coherence.
β’
Combining with Other Techniques: While min- sampling can work with other sampling methods or repetition penalties, we recommend using it as the primary truncation method to fully leverage its dynamic capabilities. Our experiments with combined samplers (see TableΒ 12) show that multiple truncation methods can lead to double normalization issues and suboptimal performance, as each methodβs hyperparameters have so far been optimised for standalone use. We discuss potential research into combined sampling in SectionΒ D.2.
Availability of Resources
To facilitate adoption, reference implementations of min- are available:
β’
Community Adoption: Min- sampling has been integrated in widely-used open-source frameworks such as Hugging Face Transformers, vLLM, SGLang and many others repositories (Zheng etΒ al., 2024), which collectively have accrued over 667,000 GitHub stars. This integration, coupled with extensive downstream usage reflected in dependents of these frameworks (e.g., over 290,000 dependent repositories for Transformers alone), underscores the methodβs availability and potential for impact (see AppendixΒ A.5 for detailed statistics). While framework popularity reflects many factors, the rapid adoption of min- demonstrates its perceived utility by developers and researchers, and its availability to the community.
β’
Project Repository: Reference code implementation is available at our project repository.111https://github.com/menhguin/minp_paper
Relevance to frontier reasoning models.
Further validating min-βs utility, Unsloth, a major inference provider whose DeepSeek R1 (DeepSeek-AI etΒ al., 2025) implementation has received > as dy 1 shows min- increasing token diversity, while Case Study 2 demonstrates min- better preserving coherence in high-confidence scenarios.
4 Case Studies: Illustrative Examples
To provide qualitative insights into how min- sampling compares to existing methods, we present two case studies that highlight differences in token truncation, especially at higher temperatures. These examples illustrate min-βs dynamic behavior of min- in practice. This visualization was originally created by open-source LLM hobbyist Maso (2024), and reproduced in this paper.
| (a) Case Study 1: Low-Certainty Next Token Prompt: "You will pay for what you have done," she hissed, her blade flashing in the moonlight. The battle that ensued ____ Token Top- Min- was 70.3 11.9 13.1 18.5 lasted 9.5 6.1 6.7 9.5 between 6.2 5.3 5.9 8.2 left 4.5 4.8 5.3 7.4 would 3.2 4.3 4.7 6.6 seemed 0.5 2.3 2.5 3.5 | (b) Case Study 2: High-Certainty Next Token Prompt: A rainbow is an optically brilliant meteorological event resulting from refraction, reflection, and dispersion of ____ Token Top- Min- light 98.3 34.4 38.2 80.9 sunlight 1.3 8.1 9.0 19.1 water 0.1 3.4 3.8 β sunshine 0.1 2.9 3.2 β a 0.05 2.7 3.0 β moisture 0.05 2.7 3.0 β |
Case Study 1: Low-Certainty Next Token
Prompt: "You will pay for what you have done," she hissed, her blade flashing in the moonlight. The battle that ensued _____
In this creative writing prompt, the model must continue a story where multiple plausible continuations exists. The next token is uncertain with a relatively flat probability distribution at high temperature.
Case Study 2: High-Certainty Next Token
Prompt: A rainbow is an optically brilliant meteorological event, resulting from the refraction, reflection, and dispersion of _____
In this factual prompt, "light" is the expected continuation, with the model highly confident in this token. We examine how various sampling methods manage this high-certainty context at .
Analysis and Insights
In other words, letβs say you are doing creative writing or fantasy roleplay. When first choosing names for new characters and settings and describing them in text, you want many diverse options (Case Study 1). Later on, when recalling these names, consistent narrative details and coherent plot points, you want high accuracy (Case Study 2).
The case studies illustrate how min- sampling dynamically adjusts the sampling threshold based on the modelβs confidence, effectively balancing creativity and coherence. In low-certainty scenarios (Case Study 1), min- behaves similarly to top-, allowing a range of plausible continuations that promote diversity without sacrificing narrative coherence. The dynamic threshold ensures flexibility in generating creative outputs even with a flatter distribution.
In high-confidence scenarios (Case Study 2), min- prioritizes the most relevant tokens, filtering out less pertinent options and maintaining factual accuracy and coherence even at high temperatures. This adaptability demonstrates min-βs ability to handle uncertain and confident contexts, ensuring robust performance by filtering out low-probability, potentially incoherent tokens.
5 Experiments
We compared min- with existing sampling methods across benchmarks and models to demonstrate its effectiveness in balancing creativity and coherence, particularly at higher temperatures.
5.1 Experimental Setup
Models
Benchmarks
We evaluate min- sampling on three diverse benchmarks:
β’
Graduate-Level Reasoning: GPQA Main Benchmark (Rein etΒ al., 2023).
β’
Grade School Math: GSM8K Chain-of-Thought (GSM8K CoT) (Cobbe etΒ al., 2021), chosen specifically because generating intermediate reasoning steps provides longer, more complex sequences that better differentiate sampling methodsβ ability to maintain coherence while introducing meaningful diversity compared to evaluating only short final answers.
β’
Creative Writing: AlpacaEval Creative Writing (Li etΒ al., 2023b).
Sampling Methods and Hyperparameters
We compared min- sampling against baseline methods, including top- sampling (Holtzman etΒ al., 2020), temperature sampling, sampling (Hewitt etΒ al., 2022b), sampling (Hewitt etΒ al., 2022b) and mirostat sampling (Basu etΒ al., 2021). Main results span temperatures 0.7 and 3.0, with further tests between 0 and 5 linked in our project repository 222https://github.com/menhguin/minp_paper/. Min- and top- perform comparably at lower temperatures 0 to 0.5 (see TableΒ 10 in Appendix), with differences within error margins. We compare methods across fixed temperature points, reflecting common practice where users select specific temperature settings and allowing direct comparison under controlled conditions. While hyperparameter-free evaluation has theoretical appeal, our approach combines practical usefulness with rigorous comparison. Nonetheless, we also separately analyze tradeoffs between accuracy and diversity across hyperparameters and note Pareto frontiers (FigureΒ 2), offering a hyperparameter-agnostic view of performance efficiency.
For min-, we used base probability thresholds of and , while top- sampling employed . These hyperparameter settings were chosen based on empirical guidelines and prior research for a fair comparison (See SectionΒ A.2 for extensive discussion).
Evaluation Metrics
Evaluation metrics were tailored to each benchmark. For GPQA and GSM8K CoT, we measured accuracy. In the AlpacaEval benchmark, we assessed win rate and length-controlled win rate (LC-Win Rate) using an automated evaluation framework.
5.2 Results
5.2.1 Graduate-Level Reasoning (GPQA Main)
Setup
The GPQA Main Benchmark consists of challenging, graduate-level multiple-choice questions in biology, physics, and chemistry. We used a 5-shot prompting strategy to provide context and improve performance, following standard practices (Rein etΒ al., 2023).
Results
TableΒ 2 presents the accuracy results on GPQA Main for different sampling methods and temperature settings using Mistral 7B. Min- sampling consistently achieves higher accuracy than others across all temperature settings. The performance gap widens at higher temperatures, demonstrating min-βs robustness in maintaining correctness while increasing diversity.
Large Model Evaluation
Evaluating min- on Mistral Large (123B parameters) shows that its advantages persist with larger models (TableΒ 3), suggesting the benefits scale with model size.
Results Across Model Families
Extensive experiments on Llama 3 (see Appendix TableΒ 8) show min-βs benefits generalize well across various model families and scales, from 1B-70B parameters.
5.2.2 Grade School Math (GSM8K CoT)
Setup
The GSM8K CoT dataset comprises 8,500 grade school math word problems (Cobbe etΒ al., 2021). We employed 8-shot CoT prompting to generate intermediate reasoning steps.
Results
TableΒ 2 shows min- outperforming other methods across most temperature settings. At lower temperatures , see TableΒ 10 in Appendix), min- and top- perform similarly, as decreasing temperature reduces the impact of all truncation methods by shrinking the token candidate pool. Min-βs advantage becomes more pronounced at higher temperatures, preserving problem-solving abilities even with increased diversity. Additionally, and sampling showed exponential runtime increases with temperature and failed completely at , while min- remained computationally efficient (SectionΒ A.3).
Notably, across all experiments, best scores were achieved with min- across a range of temperatures, rather than greedy decoding, challenging assumptions that greedy decoding is strictly optimal for task benchmarks (Castel etΒ al., 2022; Wiher etΒ al., 2022). See comparisons in TableΒ 10.
| Method | GPQA Main (5-shot) | GSM8K CoT (8-shot) | ||||||||
| Tempβ Only | 27.23 | 22.77 | 25.22 | 5.80 | 0.89 | 29.56 | 17.51 | 0.00 | 0.00 | 0.00 |
| Top- | 26.34 | 23.66 | 22.77 | 16.52 | 5.88 | 30.63 | 17.59 | 0.00 | 0.00 | 0.00 |
| Sampling | 28.13 | 25.45 | 24.55 | β | β | 32.63 | 26.99 | 0.49 | β | β |
| Sampling | 27.90 | 25.45 | 24.11 | β | β | 31.69 | 26.84 | 0.56 | β | β |
| Top- | 29.02 | 25.00 | 24.78 | 6.47 | 0.46 | 36.09 | 27.67 | 0.68 | 0.00 | 0.00 |
| Min- | 29.18 | 25.89 | 28.13 | 26.34 | 24.55 | 35.18 | 30.86 | 18.42 | 6.21 | 0.00 |
Accuracy vs. Diversity Trade-off
To further understand the accuracy-diversity tradeoff, we evaluate both metrics on GSM8K using chain-of-thought reasoning with self-consistency decoding (Wang etΒ al., 2022). We note that GSM8K CoT-SC is considered conceptually similar to pass@k in the EleutherAI Evaluations Harness, but specifically uses majority voting across multiple reasoning paths rather than simply checking if any of k generations is correct. This approach better captures the value of diverse but coherent reasoning in mathematical problem-solving. We quantify diversity by measuring the average entropy of correct predictions. Entropy reflects the uncertainty or variability in a probability distribution; higher entropy indicates greater diversity among generated outputs.
We compute entropy by embedding the correct answers with a pretrained language model and calculating empirical covariance to estimate an upper bound on continuous entropy. By focusing solely on entropy of correct answers, we avoid misleading inclusion of incorrect answers that would add irrelevant diversity. We test 20 hyperparameter combinations each for both methods.
FigureΒ 2 shows min- sampling achieves a better accuracy-creativity trade-off compared to top- sampling, consistently lying closer to the Pareto frontier and indicating more efficient performance. The greater spread of min- configurations shows its sensitivity to hyperparameter settings, allowing fine-grained control over output diversity and coherence, whereas top- configurations cluster strongly, showing less sensitivity to hyperparameter values. We further discuss this nuance in SectionΒ A.4.
5.2.3 Creative Writing
Setup
We used the AlpacaEval Creative Writing LLM-as-a-judge benchmark to assess the modelβs ability to generate creative and engaging text (Li etΒ al., 2023b); Outputs were generated by, openchat-3.5-0106, with GPT-4 Turbo as judge. Following Gusev (2023), we report both Win Rate and Length-Controlled Win Rate (LC-Win Rate). Comparison against a fixed baseline (T=1.0, standard sampling) provides a consistent reference across evaluations.
Results
TableΒ 3 shows that min- outperforms both top-, and Mirostat. Min- achieves a higher win rate, producing high-quality creative writing without sacrificing coherence.
Independent EQ-Bench Creative Writing Paech (2024) results further support these findings, with min- (, ) scoring 62 versus the baselineβs 51.5 at (Gusev, 2024). A supplementary structured LLM-as-judge evaluation focusing on specific creative writing dimensions confirmed min-βs advantages across temperature settings (see SectionΒ C.3.6).
5.3 Ablation Study
We conducted an ablation study on the AlpacaEval Creative Writing benchmark comparing min- and top- sampling across different temperatures and parameter configurations. Performance was measured using Winrate and Winrate (LC) (length-controlled).
Results in Table 7 (Appendix) show min- generally outperforming top- across the range of different temperatures and parameter settings we explored. The highest winrate is achieved with at temperature , demonstrating that min- sampling is effective in producing high-quality outputs even under conditions that promote creativity (higher temperatures). Moreover, the Winrate (LC) results are consistent with the Winrate, confirming that the min-βs benefits are robust to biases in output length differences. Confidence Intervals (95%) are shown where available.
6 Human Evaluation
To complement our quantitative benchmarks of min- sampling, we conducted a comprehensive human evaluation focusing on perceived quality and diversity of creative writing outputs across various min-, top- and temperature configurations.
Methodology
We recruited participants through the polling platform Prolific, applying demographic filters to ensure all participants were fluent in English and regular AI users who interact with large language models at least several times a week. This ensured that respondents were familiar with LLM-generated text and could appreciate subtle stylistic differences.
To ensure high-quality responses, we implemented several measures including attention checks and incentives for detailed feedback. For full details, see sectionΒ C.1. Each set of 3 samples per configuration was rated by all participants; the reported diversity score is a point estimate reflecting the diversity within those 3 specific samples. This setup aimed to provide valuable directional insights into perceived quality and diversity differences between methods, despite considerations limiting sample size such as retaining readersβ attention over very long reading durations, ensuring fair standardised scoring and keeping participant costs manageable.
1.
Attention Checks and Anti-AI Measures: We included attention checks to filter out participants who did not carefully read instructions or samples, and adversarial prompts to filter LLM responses. Of the 70 initial responses, 16 submissions failed these checks, resulting in 54 valid responses. We provide an example of our attention check below:
2.
Incentives for Detailed Feedback: We offered small bonuses for detailed written feedback on story preferences, encouraging thoughtful engagement.
| (a) Accuracy (%) on GPQA Main benchmark (Mistral Large) Method Tempβ Only 37.72 31.25 29.02 20.09 2.90 0.89 Top- 39.51 33.26 29.24 18.75 2.01 0.67 Top- 40.18 34.38 29.69 21.21 2.01 0.89 Min- (ours) 38.17 34.60 31.03 27.46 22.77 13.84 | (b) AlpacaEval Creative Writing (% win) Method Temperature Only 49.97 53.18 Mirostat 16.69 14.23 Sampling 43.50 45.51 Top- 50.43 β Min- (ours) 52.01 56.54 |
Experimental Setup
We evaluated creative writing using Llama 3 70B across different sampling configurations. The model generated stories using the prompt "Write me a creative story?" with either top- or min-. We tested three temperature settings () and two diversity levels: low (top-, min-) and high (top-, min-), yielding 12 total configurations: 2 sampling methods 3 temperatures 2 diversity settings. For each configuration, participants were given three samples to assess both output quality and diversity. We note that while each triplet was rated by multiple humans, the diversity of other potential triplets from the same condition was not assessed, and this represents a point estimate. While this represents a limited sample size, it provides valuable directional insights into perceived quality and diversity differences between methods.
Evaluation Criteria
Participants rated sets of 3 outputs from 1 (lowest) to 10 (highest) on:
1.
Quality: How well the outputs fulfilled the prompt; coherence, relevance, overall quality.
2.
Diversity: How creative or distinct the 3 stories appeared to be.
Results
TableΒ 4 summarizes average quality and diversity scores across Temperature and Diversity settings, plus a standard sampling baseline. Overall, min- sampling demonstrated advantages over top- sampling, especially at higher temperatures where top-βs quality and diversity scores often declined more sharply while min- maintained comparatively higher scores.
Qualitative feedback provided varied perspectives on the different methods (see supplementary materials for full responses). To address limitations of the initial evaluation, we conducted a refined human study with an improved inference engine and other setup refinements (detailed in SectionΒ C.2), which reinforced our findings. Both results are available on our GitHub repository333https://github.com/menhguin/minp_paper/.
These results demonstrate that min- sampling is better in both output quality and diversity.
| Lower Diversity Settings | Higher Diversity Settings | ||||||
| Temp | Metric | Standard | Top- | Min- | Standard | Top- | Min- |
| Sampling | = 0.1 | = 0.2 | Sampling | = 0.9 | = 0.05 | ||
| T=1 | Quality | 7.550.2365 | 5.960.3077 | 7.060.2039 | 7.960.1532 | 7.670.1931 | 8.020.1858 |
| Diversity | 7.080.2921 | 2.400.2765 | 5.830.2795 | 7.660.2014 | 7.040.2587 | 7.740.2235 | |
| T=2 | Quality | 7.760.2640 | 5.430.3083 | 7.620.2105 | 7.850.1931 | 7.750.1885 | 7.980.1952 |
| Diversity | 7.660.2459 | 1.830.2212 | 6.910.2658 | 7.570.2058 | 7.660.2066 | 7.960.2117 | |
| T=3 | Quality | 6.750.3509 | 5.750.3201 | 7.740.2418 | 6.830.2895 | 7.110.2869 | 7.570.2303 |
| Diversity | 7.040.2836 | 2.250.3353 | 7.600.2550 | 7.430.2830 | 7.490.2396 | 7.660.1996 | |
7 Conclusion
In this paper, we introduced min- sampling, a novel truncation method for LLMs that dynamically adjusts sampling thresholds based on model confidence. Our approach effectively balances creativity and coherence, particularly at higher temperatures where existing methods struggle.
Through comprehensive experiments across diverse benchmarks, we demonstrate that min- sampling consistently outperforms existing methods in both quality and diversity. Human evaluations confirm these advantages, showing a clear preference for min- outputs in real-world applications.
Min-βs strengths lie in its simplicity, computational efficiency, and seamless integration into existing pipelines, as evidenced by its rapid adoption in many leading open-source frameworks. By addressing the longstanding diversity-quality tradeoff in a way practitioners have found immediately useful, min- represents a notable advancement in generative language modeling, enhancing applications that require both high-quality and diverse text generation. For more detailed examples, see sectionΒ D.1.
Reproducibility Statement
We have made demonstrable efforts to ensure the reproducibility of our results. The implementation of the proposed min- sampling method is provided in SectionΒ A.1 and is also available at our project repository.444https://github.com/menhguin/minp_paper/ Detailed descriptions of experimental setups, including model configurations, hyperparameter settings, and evaluation protocols, are outlined in SectionΒ 5 and SectionΒ A.2. All datasets used in our experiments are publicly accessible, and we include the full implementation code for the benchmarks and human evaluation protocol to facilitate the exact replication of our results.
Additionally, due to min-βs popularity within the open-source community, we note that there are multiple, unaffiliated and independent replications of the benchmarks and claims presented in this paper. Most notably, our claims of min- improving medium-to-high temperature GPQA and GSM8K evaluation results have been replicated by Top-N sampling developers (Tang etΒ al., 2024), and independent AlpacaEval Creative Writing and EQ-Bench evaluations were conducted by Ilya Gusev (Gusev, 2023; 2024), with results broadly confirming our findings about min-βs advantages.
Ethics Statement
Min- sampling aims to improve the diversity and coherence of text generated by large language models. We acknowledge the following ethical considerations:
β’
Potential misuse: Min- could potentially enhance the fluency of misleading or harmful content. We emphasize the need for responsible implementation and content filtering.
β’
Safety risks: It is possible that high-temperature text generation increases risks of circumventing safety finetuning, although, in practice, we are not aware of such instances.
β’
Transparency: To ensure reproducibility and further research, we have open-sourced our implementation and provided extensive details on the experimental setup and results. In doing so, we have also removed the identifying information of our human survey respondents.
β’
AI Safety and Interpretability: We are actively exploring follow-up work applying min- to mechanistic interpretability for AI safety and alignment. Specifically, we are investigating its use in uncertainty-aware generation, neuron activation filtering, and structured latent selection to enhance model robustness and truthfulness. We discuss an extension of this approach in Β§D.2, where we introduce min-, a variation that leverages standard deviation-based filtering for broader applications.
Acknowledgements
Min- began as a passion project to see if small, open-source language models could write diverse and interesting stories. As such, its creation is a winding story with a diverse set of characters.
We thank Apart Research, particularly Esben Kran and Jason Schreiber, for providing extensive guidance throughout our Apart Lab Fellowship, during which min- evolved from an interesting idea to a full, empirically validated research paper. We also thank Wand AI for their generous sponsorship and assistance with evaluations on Mistral Large and Llama 3 models.
Our gratitude extends to the numerous members of the open-source LLM community who supported the research and adoption of min-. We are particularly indebted to Ilya Gusev for his independent evaluations, Romain Dal Maso for his intuitive visualizations, Hailey Schoelkopf for her extensive technical support, and JoΓ£o Gante for his help with the Hugging Face Transformers library.
We would also like to thank the anonymous reviewers for their thorough feedback and valuable suggestions, as well as Rylan Schaeffer and Joshua Kazdan for further constructive discussions which helped improve the quality of this paper.
While our coauthors are employed in various research roles, min- does not belong to any institution. Our technical contributions and research findings are free for use by the open-source community.
Contribution Statement
Andrew Baker conceived the original idea of min- sampling, tested and shared it with hobbyists within the open-source LLM community, leading to its rapid and widespread external adoption.
Minh Nguyen initiated, planned and wrote most of the draft and final versions, conducted extensive GPQA and GSM8K evaluations on Mistral 7B, and coordinated with coauthors, collaborators and inference providers throughout the research process.
Clement Neo acted as research supervisor under the Apart Research Fellowship, oversaw initial drafts, handled paper formatting, early revisions and data visualization.
Andreas Kirsch conducted GSM8K Self-Consistency evaluations to assess Diversity-Accuracy tradeoffs and helped flesh out theoretical explanations of min-.
Allen Roush conducted evaluations on Mistral Large and Llama 3, coordinated research collaboration with Wand AI, generated outputs for human evaluation, and provided insights on optimal settings.
Ravid Shwartz-Ziv rewrote a large portion of the final draft, proposed the inclusion of a human evaluation, and provided guidance on evaluations, data presentation and academic publishing standards.
Appendix A Implementation Details
A.1 Min- Implementation (code listing)
Below is the implementation code for min- truncation sampling as detailed in the Hugging Face Transformers library, with range exception handling and keeping minimum tokens to prevent errors.
This implementation code, along with other similar implementations in other open-source inference engines, logs of automated evaluations for GPQA, GSM8K Chain-of-Thought and AlpacaEval Creative Writing is available at https://github.com/menhguin/minp_paper/.
A.2 Hyperparameters Settings (Selection Methodology)
To choose fair and optimal hyperparameter settings, we mainly cross-referenced publicly-reported scores on MAUVE (Pillutla etΒ al., 2023), common recommendations from leading model providers, and any recommendations from the original authors. We also found that Risk Levels (Zhou etΒ al., 2024) correlate strongly with optimal results across temperature ranges. Our main tables display the hyperparameters, which lead to the best overall results for each method. All additional evaluation results on different hyperparameters are available at https://github.com/menhguin/minp_paper/
For min-, base probability thresholds of and were used, while top- sampling employed and .
For min- = 0.05 and 0.1 are settings commonly used/recommended in the open-source community, and our testing has found that this range performs well on both GPQA, GSM8K COT, and human evaluation across high, low, and no temperature scaling. We also tested min- = 0.2 and min- = 0.3, but these are not commonly used.
For top-, top- = 0.9 and top- =0.95 are settings commonly used/recommended in the open-source community, and found in several independent MAUVE assessments to be optimal (Hewitt etΒ al., 2022a; Zhu etΒ al., 2024). We mainly reference the Risk Levels framework from Zhou etΒ al. (2024), which measures tradeoffs between diversity and risk/precision in text generation, specifically Risk Level 15 for Mixtral 7B, which we used as a reference point for the top-, and sampling settings.
| Model | Method | Parameter | Risk Std Error β | Recall β |
| Mixtral-7b | Top-k | 181 | 1.759 | 0.364 |
| Top- | 0.9315 | 6.315 | 0.447 | |
| Adaptive | 2.2e-5 | 2.757 | 0.466 | |
| Eta | 1.96e-4 | 4.712 | 0.505 | |
| Mirostat | 6.71 | 2.213 | 0.468 |
For top-, we could not find clear recommendations on the optimal hyperparameters. We conducted tests on k = 10, 15, 20, 40, 50 and 180. Due to the nature of top-, we noted that best scores and settings varied substantially by temperature, making a fair comparison difficult as, in practice, top- is meant to be a static threshold and not dynamically adjusted at inference. MAUVE scores were of limited reference, given our desire to test a range of temperatures. Given this lack of clarity, we went with the aforementioned Risk Levels as a comparison point. (Zhou etΒ al., 2024)
For and , we found inter-agreement between the authorβs original recommendation, independent MAUVE assessments (Zhu etΒ al., 2024), and Risk Levels. We tested and values 0.0002 and 0.0009, found 0.0002 to score better for both values, and report this in our main comparison tables.
A.3 Test Time Compute challenges
While running GPQA and GSM8K CoT for and sampling via Hugging Face and the EleutherAI Evaluation Harness (Gao etΒ al., 2023), we noted that test-time compute increased exponentially with temperature. During our experiments, min-, top-, and top- generally took 2-5 minutes to evaluate on Mistral 7B at every temperature for both GPQA and GSM8K. Runtime on an A100 Colab increased from Β 5 minutes at to 8 minutes at and 30 minutes at . Neither nor functioned at . On GSM8K, and each took 2 hours to evaluate, equal to the average for our Llama 70B/Mistral Large run. This time also scaled exponentially with increased temperature.
Our experience suggests min- is a practical alternative to and samplingβs entropy-based heuristics both on quantitative benchmarks and compute efficiency.
A.4 How percentages thresholds differ for Min-p and Top-p
In choosing hyperparameter values, min- and top-βs percentage thresholds differ in subtle but meaningful ways. Strictly speaking, min-βs threshold is not the same as an equivalent "1 - top-" threshold. For example, when top- = 0.9, the last <10% of the total distribution is truncated. However, itβs possible for min- = 0.1 to truncate more than 10% of a distribution.
Consider the following top 5 token probabilities: 80%, 7%, 3%, 2%, 1%. With top- set to 0.9, the top 3 tokens comprising 90% of the distribution is preserved. With min- = 0.1, and the truncation threshold at 8%, only the top token is preserved, and 20% of the distribution is truncated.
In practice, this means that in high-certainty token distributions, min- truncates a larger percentage of that probability distribution than its value. This contributes to min-βs ability to consistently choose high-certainty tokens despite high temperature scaling.
Hence, low values (such as from 0.01 to 0.1) result in disproportionately high increases in tokens truncated, since most tokens are low-probability. This results in Figure Β 2βs observation that min-βs is more sensitive than top-βs when adjusted by the same percentage values/basis points.
A.5 Detailed Community Adoption Statistics
To provide further transparency regarding the adoption of Min- sampling, TableΒ 6 summarizes key statistics from major projects that have integrated the method.
| Repository | Stars | Downstream Repos |
| Hugging Face Transformers | 144,000 | 291,000 |
| Ollama | 141,000 | 400 |
| open-webui | 95,100 | β |
| llama.cpp | 80,400 | β |
| vLLM | 47,500 | β |
| text-generation-webui | 43,600 | β |
| unsloth | 38,900 | β |
| continue | 26,300 | β |
| SGLang | 14,400 | β |
| sillytavern | 14,500 | β |
| mamba | 14,900 | β |
| kobold.cpp | 7,400 | β |
| Total | 668,000 | 291,400+ |
Clarification on Community Adoption Metrics:
We wish to clarify that our earlier claim specifying "54,000 repositories and 1.1 million stars" was based on preliminary GitHub scans using search terms like "min_p" which produced many false positives. An independent contributor has discussed with us the limitations of this search approach Nguyen (2025).
For this revised assessment, we focused on identifying verified integrations in high-traffic frameworks where min- is explicitly implemented as a sampling option in the official codebase. This approach, while potentially undercounting total adoption, provides a more conservative and verifiable measure of min-βs integration in the open-source ecosystem.
Given the difficulty of obtaining direct usage statistics, as many leading providers do not specifically collect or disclose such data (per our enquiries with Hugging Face and LMSys staff), we believe this evidence of integration into major frameworks represents the most reliable indicator of min-βs practical availability and potential for impact.
Supplementary Search Evidence. Using the following GitHub code search query:
This slightly restrictive search, which assumes min- is a variable present in a file alongside several other sampling parameters, yields over 13,000 matching code files. A manual scan suggests that many of these references are indeed related to sampling configurations, corroborating the methodβs broader community-level diffusion beyond the curated list above.
Appendix B Benchmark Evaluation Results
B.1 Ablation Study
The results in Table 7 show that min- sampling generally outperforms top- sampling across different temperatures and parameter settings. The highest winrate is achieved with at temperature , demonstrating that min- sampling is effective in producing high-quality outputs even under conditions that promote creativity (higher temperatures). Moreover, the Winrate (LC) results are consistent with the Winrate results, confirming that the benefits of min- sampling are robust to biases due to differences in output length. Confidence intervals (95%) are shown where available. This table includes results from exploring a range of hyperparameters for each method.
| Model | Winrate (%) 95% CI | Winrate (LC) (%) 95% CI | Avg Length (# runs) |
| temp150_minp_10 | 56.54 3.68 | 58.12 3.67 | 1852 (696) |
| temp150_minp_15 | 53.37 3.70 | 56.73 3.67 | 1816 (698) |
| temp150_minp_20 | 53.38 3.71 | 55.45 3.69 | 1835 (696) |
| quad_20_100 | 52.80 3.71 | 55.43 3.69 | 1821 (696) |
| temp100_minp_05 | 52.01 3.71 | 55.07 3.69 | 1808 (697) |
| temp200_minp_20 | 53.08 3.70 | 54.82 3.69 | 1861 (698) |
| temp80_topp_98 | 51.29 3.71 | 54.65 3.69 | 1810 (697) |
| dynatemp_50_150_75_minp_05 | 51.58 3.72 | 54.42 3.69 | 1807 (695) |
| dynatemp_50_200_100_minp_10 | 51.87 3.71 | 54.33 3.69 | 1825 (696) |
| temp150_minp_10_seed1337 | 52.86 3.70 | 53.84 3.69 | 1856 (699) |
| temp170_minp_15 | 52.65 3.70 | 53.75 3.69 | 1855 (697) |
| temp120_minp_10 | 51.36 3.71 | 53.75 3.69 | 1829 (698) |
| quad_15_100 | 51.65 3.71 | 53.70 3.69 | 1843 (697) |
| tfs_95 | 50.79 3.71 | 53.49 3.69 | 1802 (696) |
| tfs_98 | 50.72 3.71 | 53.39 3.69 | 1807 (697) |
| temp100_minp_10 | 50.14 3.72 | 53.24 3.69 | 1793 (695) |
| temp100_topp_98 | 50.43 3.71 | 53.00 3.69 | 1834 (697) |
| temp100_topp_90 | 50.07 3.73 | 52.57 3.70 | 1815 (692) |
| temp80 | 49.28 3.71 | 52.40 3.69 | 1797 (698) |
| temp100_topp_95 | 50.22 3.71 | 51.80 3.69 | 1835 (696) |
| temp100_minp_02 | 50.43 3.72 | 51.62 3.69 | 1853 (696) |
| temp80_minp_02 | 48.85 3.71 | 51.46 3.69 | 1802 (696) |
| temp80_minp_05 | 47.84 3.71 | 50.99 3.69 | 1808 (696) |
| temp80_topp_95 | 48.78 3.71 | 50.76 3.69 | 1793 (696) |
| temp100 | 50.00 | 50.00 | 1902 (700) |
| dynatemp_100_250_100_minp_10 | 50.86 3.71 | 50.00 3.69 | 2227 (698) |
| quad_25_100 | 47.85 3.71 | 49.94 3.69 | 1807 (697) |
| temp150_tfs_95 | 51.08 3.71 | 49.94 3.69 | 1969 (697) |
| greedy | 46.64 3.70 | 49.90 3.69 | 1765 (700) |
| temp150_minp_05 | 48.57 3.71 | 48.13 3.69 | 1919 (697) |
| temp150_topp_80 | 24.21 4.42 | 17.65 3.91 | 5258 (95) |
| temp150_minp_02 | 44.83 3.70 | 42.09 3.67 | 2149 (696) |
| dynatemp_50_150_100 | 35.37 3.55 | 34.94 3.52 | 2764 (697) |
| mirostat_40_10 | 16.69 2.76 | 16.04 2.72 | 1848 (698) |
| mirostat_50_10 | 16.40 2.75 | 16.04 2.72 | 1822 (698) |
| mirostat_60_10 | 15.62 2.69 | 14.97 2.64 | 1838 (698) |
| temp150_topp_98 | 0.00 | 0.00 0.03 | 3125 (97) |
| temp150_topp_95 | 1.03 1.03 | 0.61 0.79 | 4665 (97) |
| temp150_topp_90 | 2.06 1.45 | 0.93 0.98 | 6139 (97) |
Evaluations were performed using GPT-4 Turbo with chain-of-thought prompting on outputs from the openchat-3.5-0106 model. Note: Throughout the rest of our paper, we choose to compare Min- = 0.05 or 0.1 vs Top- = 0.9 or 0.95. However, since the original implementation of this evaluation includes additional values, we now also include this.
B.2 Results of GPQA Main or GSM8K CoT benchmarks for Llama 3 models
| Temperature | 0.0 | 0.3 | 0.5 | 0.7 | 1.0 | 1.5 | 2.0 | 3.0 | 4.0 | 5.0 |
| Tempβ Only | 28.57 | 28.57 | 25.89 | 26.79 | 24.55 | 12.95 | 3.35 | 2.46 | 2.23 | 2.01 |
| Top- = 0.9 | 28.57 | 27.23 | 28.79 | 27.23 | 25.45 | 18.75 | 3.57 | 2.68 | 2.01 | 2.23 |
| Top- = 0.95 | 28.57 | 29.24 | 26.34 | 26.56 | 25.67 | 19.64 | 6.03 | 2.68 | 2.46 | 2.90 |
| Min- = 0.05 | 28.57 | 29.46 | 30.13 | 27.46 | 23.88 | 23.44 | 22.32 | 16.52 | 6.47 | 3.12 |
| Min- = 0.1 | 28.57 | 27.46 | 28.57 | 27.01 | 26.56 | 25.67 | 21.43 | 19.42 | 14.29 | 6.92 |
| Temperature | 0.0 | 0.3 | 0.5 | 0.7 | 1.0 | 1.5 | 2.0 | 3.0 | 4.0 | 5.0 |
| Tempβ Only | 46.55 | 45.03 | 42.99 | 37.60 | 28.89 | 0.23 | 0.00 | 0.00 | 0.00 | 0.00 |
| Top- = 0.9 | 46.55 | 45.19 | 44.20 | 40.11 | 37.23 | 5.23 | 0.00 | 0.00 | 0.00 | 0.00 |
| Top- = 0.95 | 46.55 | 44.73 | 44.28 | 41.62 | 33.89 | 2.50 | 0.00 | 0.00 | 0.00 | 0.00 |
| Min- = 0.05 | 46.55 | 43.67 | 45.11 | 42.23 | 36.24 | 24.64 | 7.05 | 0.00 | 0.00 | 0.00 |
| Min- = 0.1 | 46.55 | 45.49 | 43.63 | 42.68 | 40.18 | 29.11 | 17.06 | 9.82 | 0.15 | 0.00 |
| Temperature | 0.0 | 0.3 | 0.5 | 0.7 | 1.0 | 1.5 | 2.0 | 3.0 | 4.0 | 5.0 |
| Tempβ Only | 27.23 | 25.89 | 24.55 | 27.68 | 25.00 | 20.09 | 5.36 | 2.23 | 2.23 | 1.79 |
| Top- = 0.9 | 27.23 | 29.46 | 27.68 | 28.79 | 30.36 | 25.45 | 9.82 | 2.68 | 2.68 | 1.79 |
| Top- = 0.95 | 27.23 | 28.79 | 27.23 | 27.68 | 29.46 | 23.00 | 5.58 | 3.13 | 1.79 | 1.79 |
| Min- = 0.05 | 27.23 | 28.35 | 27.46 | 27.23 | 32.37 | 27.68 | 27.46 | 21.65 | 11.38 | 3.79 |
| Min- = 0.1 | 27.23 | 28.35 | 28.79 | 31.70 | 29.24 | 31.25 | 23.66 | 23.66 | 16.96 | 8.93 |
| Temperature | 0.0 | 0.3 | 0.5 | 0.7 | 1.0 | 1.5 | 2.0 | 3.0 | 4.0 | 5.0 |
| Tempβ Only | 76.72 | 77.10 | 76.42 | 74.00 | 64.59 | 3.11 | 0.00 | 0.00 | 0.00 | 0.00 |
| Top- = 0.9 | 76.72 | 76.65 | 76.72 | 75.66 | 73.31 | 28.81 | 0.00 | 0.00 | 0.00 | 0.00 |
| Top- = 0.95 | 76.72 | 77.41 | 77.63 | 76.50 | 71.11 | 16.83 | 0.00 | 0.00 | 0.00 | 0.00 |
| Min- = 0.05 | 76.72 | 76.12 | 76.50 | 75.51 | 73.24 | 57.92 | 26.61 | 0.15 | 0.00 | 0.00 |
| Min- = 0.1 | 76.72 | 77.18 | 75.51 | 75.36 | 73.01 | 75.44 | 52.08 | 2.50 | 0.00 | 0.00 |
| Temperature | 0.0 | 0.3 | 0.5 | 0.7 | 1.0 | 1.5 | 2.0 | 3.0 | 4.0 | 5.0 |
| Tempβ Only | 29.02 | 27.46 | 28.79 | 29.91 | 30.36 | 22.99 | 6.92 | 3.12 | 2.46 | 3.35 |
| Top- = 0.9 | 29.02 | 28.57 | 29.69 | 30.80 | 28.79 | 24.55 | 9.38 | 2.46 | 2.46 | 2.90 |
| Top- = 0.95 | 29.02 | 30.36 | 31.92 | 27.68 | 30.58 | 25.67 | 10.04 | 2.68 | 2.90 | 2.90 |
| Min- = 0.05 | 29.02 | 30.13 | 31.70 | 30.80 | 28.35 | 26.34 | 29.24 | 22.77 | 9.82 | 5.13 |
| Min- = 0.1 | 29.02 | 30.13 | 29.69 | 29.24 | 29.24 | 32.14 | 25.22 | 22.99 | 20.54 | 12.50 |
| Temperature | 0.0 | 0.3 | 0.5 | 0.7 | 1.0 | 1.5 | 2.0 | 3.0 | 4.0 | 5.0 |
| Tempβ Only | 84.91 | 84.61 | 84.84 | 81.50 | 75.21 | 10.39 | 0.00 | 0.00 | 0.00 | 0.00 |
| Top- = 0.9 | 84.91 | 84.91 | 84.08 | 83.24 | 80.36 | 48.37 | 0.08 | 0.00 | 0.00 | 0.00 |
| Top- = 0.95 | 84.91 | 84.23 | 84.08 | 82.26 | 80.06 | 32.52 | 0.00 | 0.00 | 0.00 | 0.00 |
| Min- = 0.05 | 84.91 | 85.06 | 84.31 | 83.32 | 80.44 | 67.25 | 32.15 | 0.08 | 0.00 | 0.00 |
| Min- = 0.1 | 84.91 | 84.46 | 84.84 | 82.87 | 82.18 | 75.44 | 52.08 | 2.50 | 0.00 | 0.00 |
| Temperature | 0.5 | 0.7 | 1.0 | 2.0 | 3.0 |
| Tempβ Only | 40.85 | 39.51 | 41.07 | 5.58 | 2.46 |
| Top- = 0.9 | 40.85 | 42.19 | 40.63 | 7.81 | 3.35 |
| Top- = 0.95 | 40.63 | 41.29 | 39.51 | 6.47 | 2.68 |
| Min- = 0.05 | 40.85 | 41.52 | 38.62 | 33.71 | 23.88 |
| Min- = 0.1 | 41.07 | 42.19 | 41.52 | 33.04 | 24.55 |
| Min- = 0.2 | 43.30 | 41.96 | 40.40 | 34.38 | 31.47 |
| Temperature | 0.7 | 3.0 |
| Tempβ Only | 93.33 | 0.08 |
| Top- = 0.9 | 93.48 | 0.08 |
| Min- = 0.05 | 93.03 | 6.07 |
| Min- = 0.2 | 92.42 | 61.03 |
B.3 Results of GPQA Main or GSM8K CoT Benchmarks for Mistral 7B Models - Low Temperatures (T = 0.5)
| Temperature | 0.0 | 0.05 | 0.1 | 0.2 | 0.3 | 0.5 |
| Temp Only | 27.68 | 27.68 | 26.34 | 25.22 | 24.33 | 24.11 |
| Top- = 0.9 | 27.68 | 28.13 | 27.23 | 26.56 | 24.78 | 24.55 |
| Top- = 0.95 | 27.68 | 27.90 | 27.23 | 25.22 | 24.55 | 24.11 |
| Min- = 0.05 | 27.68 | 27.90 | 27.23 | 25.45 | 24.55 | 24.33 |
| Min- = 0.1 | 27.68 | 28.35 | 27.46 | 26.56 | 24.33 | 24.33 |
| Temperature | 0.0 | 0.05 | 0.1 | 0.2 | 0.3 | 0.5 |
| Temp Only | 39.35 | 38.59 | 38.21 | 38.59 | 37.23 | 36.32 |
| Top- = 0.9 | 39.35 | 39.27 | 40.03 | 39.27 | 37.53 | 38.36 |
| Top- = 0.95 | 39.35 | 38.74 | 38.74 | 39.58 | 37.38 | 36.62 |
| Min- = 0.05 | 39.35 | 38.59 | 39.65 | 40.33 | 38.44 | 37.68 |
| Min- = 0.1 | 39.35 | 39.20 | 38.51 | 38.21 | 38.06 | 37.07 |
B.4 Results of GPQA Main or GSM8K CoT Benchmarks for Mistral 7B Models - Combined Top P and Top K sampling
Top- = 0.5
Top- = 0.9
Top- = 0.95
Top- = 1.0
B.5 Results of High Minimum Probability Sampling with Mistral-7B on Mathematical Reasoning Tasks
| Temperature | min_p | GSM8K Score | GPQA Score |
| 3.0 | 0.7 | 0.4041 | 0.2478 |
| 0.6 | 0.3533 | 0.2500 | |
| 0.5 | 0.3237 | 0.2433 | |
| 0.4 | 0.2297 | 0.2299 | |
| 0.3 | 0.1327 | 0.2746 | |
| 2.0 | 0.7 | 0.4071 | 0.2455 |
| 0.6 | 0.4162 | 0.2567 | |
| 0.5 | 0.3738 | 0.2567 | |
| 0.4 | 0.3397 | 0.2545 | |
| 0.3 | 0.3078 | 0.2813 | |
| 1.0 | 0.7 | 0.4215 | 0.2790 |
| 0.6 | 0.3958 | 0.2545 | |
| 0.5 | 0.4124 | 0.2478 | |
| 0.4 | 0.4185 | 0.2589 | |
| 0.3 | 0.3942 | 0.2522 | |
| 0.7 | 0.7 | 0.4412 | 0.2835 |
| 0.6 | 0.4208 | 0.2835 | |
| 0.5 | 0.4177 | 0.2813 | |
| 0.4 | 0.4200 | 0.2567 | |
| 0.3 | 0.4155 | 0.2679 | |
| 0.5 | 0.7 | 0.4200 | 0.2857 |
| 0.6 | 0.4155 | 0.2835 | |
| 0.5 | 0.4147 | 0.2835 | |
| 0.4 | 0.4117 | 0.2746 | |
| 0.3 | 0.4359 | 0.2612 |
B.6 Greedy Decoding Model Performance on GPQA and GSM8K
| Model | Greedy Score (T=0) | Best Score | Hyperparameters | Temperature |
| GPQA (<10B Models) | ||||
| Mistral 7B | 27.35% | 29.18% (+1.83%) | Min P = 0.1 | 0.7 |
| Llama 3.2 3B | 27.23% | 32.37% (+5.14%) | Min P = 0.05 | 1.0 |
| Llama 3.1 8B | 29.02% | 32.15% (+3.13%) | Min P = 0.1 | 1.5 |
| GPQA (Larger Models) | ||||
| Llama 3.1 70B | 41.07% | 43.30% (+2.23%) | Min P = 0.2 | 0.3 |
| Llama 3.1 70B | 41.07% | 42.19% (+1.12%) | Min P = 0.1 | 0.5 |
| Llama 3.1 70B | 41.07% | 41.52% (+0.45%) | Min P = 0.1 | 1.0 |
| GSM8K | ||||
| Mistral 7B | 39.35% | 40.33% (+0.98%) | Min P = 0.05 | 0.3 |
| Llama 3.1 8B | 84.91% | 85.06% (+0.15%) | Min P = 0.05 | 0.2 |
Appendix C Additional Evaluations
C.1 Original Human Evaluation Survey Methodology
Survey Links:
β’
Survey: https://forms.gle/WUXPnSWkZq6uScbz9
β’
Results: Available in our linked Github repository.
C.1.1 Survey Implementation Details
1. Participant Recruitment
β’
Platform: Prolific Academic
β’
Sample Size: Initial n=70, Final n=54 after attention check filtering
β’
Demographic Requirements:
β
Fluent English speakers
β
Regular AI users (self-reported interaction with LLMs at least several times per week)
β
18+ years old
β
No technical AI/ML knowledge required
β
2. Recruitment Notice
Participants were recruited with the following study description:
Title: βIs our new AI model better at Creative Writing?β
Background provided to participants:
βIn this study, you will evaluate AI-generated text from Large Language Models (LLMs), which are AI systems designed to generate human-like text (e.g. ChatGPT). Weβre investigating different methods of generating text from these models and how humans perceive the quality and diversity of the outputs.
We are testing the creative writing prompt: βWrite me a creative story?ββ
3. Survey Structure
β’
Format: Google Forms
β’
Duration: Average completion time 25-30 minutes
β’
Compensation: Base rate Β£6.00 (Β£12/hour) with potential Β£1.00+ bonus for detailed qualitative feedback
β’
Question Types: Mix of scale ratings (1-10) and open-ended responses
Story outputs were generated using Llama 3 70B with consistent prompt (βWrite me a creative story?β) across all conditions. The survey consisted of 6 sections evaluating different temperature/diversity settings:
β’
A. Temperature 1.0 - Low Diversity (min p = 0.2, top p = 0.1)
β’
B. Temperature 2.0 - Low Diversity
β’
C. Temperature 3.0 - Low Diversity
β’
D. Temperature 1.0 - High Diversity (min p = 0.05, top p = 0.9)
β’
E. Temperature 2.0 - High Diversity
β’
F. Temperature 3.0 - High Diversity
For each section, participants evaluated:
β’
3 outputs from Model A (control/baseline/pure sampling/standard sampling/no truncation sampling)
β’
3 outputs from Model B (top- sampling)
β’
3 outputs from Model C (min- sampling)
4. Rating Criteria
For each model in each section, participants provided two ratings:
Quality Rating (1-10 scale)
β’
Prompt: βModel [X] output quality (how well you think it fulfilled the prompt)β
β’
Scale anchors: 1=βHorribleβ to 10=βAmazingβ
Diversity Rating (1-10 scale)
β’
Prompt: βModel [X] output diversity (how different/distinct the 3 stories feel from each other)β
β’
Scale anchors: 1=βUninterestingβ to 10=βDiverseβ
5. Sample Presentation
β’
Model outputs were presented in consistent formatting
β’
Each output clearly labeled (Sample 1, Sample 2, Sample 3)
β’
Models were anonymized as βModel A/B/Cβ
β’
All participants saw identical story outputs
Example Story Outputs
Below are representative samples from one setting (Temperature 1.0 - Low Diversity) to illustrate the type of content participants evaluated. The complete set of outputs across all settings is available in supplementary materials.
6. Quality Control Measures
Attention Checks
Two strategically placed attention checks were embedded in the survey at the 40%Β and 70%Β marks:
First Check (embedded inside Temperature Setting B samples):
Second Check (embedded inside Temperature Setting D samples):
These checks served multiple purposes:
β’
Verified careful reading of instructions
β’
Detected automated/LLM responses
β’
Ensured engagement throughout the survey
β’
Required basic memory of story content
Participants who failed to follow these instructions in their final submission (n=16) were excluded from analysis.
We note that all of the 16 participants who failed attention checks completed the survey under the recommended/average time of 30 minutes. Rejected submissions took on average Β 15 minutes. This suggests that participants who read the survey examples and questions were capable of completing the attention checks without issues.
Engagement Validation
β’
Required minimum 1-2 sentence explanation for model preferences
β’
Offered bonus incentive for detailed qualitative feedback. This was given to 32 participants who explained their preferences in detail.
β’
Manual review of open-ended responses for signs of low effort/automated completion. 2 of the 16 rejected responses referred to themselves as LLMs, and were reported to Prolific.
Response Time Monitoring
β’
Tracked total completion time
β’
Flagged suspiciously quick completions (15 minutes) for manual review, cross referenced with response quality and attention check completion
7. Open-Ended Questions
1.
Model Preference: βWhich Model(s) on which Settings did you like the most overall? What did you like about it? Please explain in at least 1-2 sentences.β
2.
Comparison to Known AI: βWhich AI chatbots do you regularly use, if any (e.g. ChatGPT, Claude, Gemini)? If so, how well did the best Model here perform in creative writing, compared to what youβve used?β
3.
Additional Comments: βAny other comments/anything that stood out to you?β
8. Data Collection & Processing
β’
Responses collected via Google Forms
β’
Raw data exported to CSV for analysis (available on Github repo)
β’
Quality control filtering applied before analysis. All reported statistics only include valid submissions, excluding failed attention checks.
β’
Statistical analysis performed using paired t-tests and inter-annotator agreement
β’
Qualitative responses coded for common themes
C.2 Additional Human Evaluation with VLLM Inference Engine
To address limitations in our initial human evaluation, we conducted a second evaluation with several key methodological improvements. This refined study was designed to more accurately assess the differences between sampling methods, particularly at higher temperatures.
Inference Engine Change.
The primary methodological change was switching from Hugging Faceβs Transformers library to VLLM for text generation. This change was critical because we discovered that Hugging Face applies temperature scaling after truncation sampling (rather than before), which considerably reduces the effect of truncation sampling methods, especially at higher temperatures.
The VLLM inference engine correctly applies temperature scaling before truncation, allowing us to properly evaluate the intended behavior of min- and top- sampling as described in our methodology. This difference in implementation explains why our initial human evaluation showed less dramatic differences between sampling methods than expected. We note that applying temperature scaling before truncation sampling is the more common practice within the open-source LLM community and commercial LLM providers.
Methodological Improvements.
Beyond the inference engine change, we implemented several other improvements:
β’
Higher-quality participant pool: Utilized Prolificβs newly introduced "AI Testers" feature, providing a vetted pool of respondents with experience judging generative AI outputs
β’
Standard hyperparameter selection: Compared commonly-used settings: top- = 0.9 and 0.95 with min- = 0.05 and 0.1
β’
Full-length story outputs: Replaced 3 short paragraph samples per setting with a single, complete story. This better represents a full LLM response to the writing prompt and emphasizes meaningful differences in longform textual coherence/degradation
β’
Increased engagement time: Extended allotted reading time from 30 to 45 minutes and increased compensation from Β£6.00 to Β£9.00 to account for longer stories
β’
Enhanced attention checks: Updated checks to detect LLM-assisted responses from the latest ChatGPT, Claude and Gemini models
β’
Clearer evaluation criteria: Added explicit instructions for participants to read closely and penalize narrative incoherence, as previous participants tended to overlook incoherent text
Example of Incoherence Overlooked in Initial Study.
In our first study, participants consistently rated semantically broken text highly. For instance, this Standard Sampling excerpt (at temperature 3.0) received a 7-8/10 rating despite clear textual incoherence:
For reference, when we asked leading commercial LLMs (ChatGPT and Claude) to rate this standard sampling passage, they scored it between 3-4/10 due to grammatical errors and narrative incoherence, while rating top- 5-6.5/10 and min-βs more coherent excerpts around 7-8.5/10.555These chat logs can be viewed at: Claude Sonnet 3.7: https://claude.ai/share/79230c92-3b5b-40d6-9d2d-1f59d3061fa8 and GPT 4.5: https://chatgpt.com/share/67d5540a-8bd4-8010-822d-a28c58a7d740
Results and Implications.
As shown in TableΒ 15, this refined evaluation reveals dramatic differences between sampling methods at higher temperatures. At temperature 3.0, standard sampling and top- both produced incoherent text (scores near 1 out of 10), while min- maintained notably higher quality (5.80 vs 1.23) and diversity (5.90 vs 1.27) scores. This striking performance gap validates min-βs core advantage: maintaining coherence at higher temperatures where other methods falter.
This advantage is particularly evident under high-temperature conditions where baseline methods degrade significantly, allowing min- to maintain relative coherence in these regimes, even if absolute quality scores are lower than those achievable in optimal low-temperature settings.
Methodological Philosophy.
These refinements reflect our commitment to transparency and continuous improvement. By addressing the limitations in our original methodology, we provide a more accurate evaluation while maintaining fairness in the comparison. Both evaluation results are included to give a comprehensive view of min-βs performance advantages.
| Lower Diversity Settings | Higher Diversity Settings | ||||||
| Temp | Metric | Standard | Top- | Min- | Standard | Top- | Min- |
| Sampling | = 0.9 | = 0.1 | Sampling | = 0.95 | = 0.05 | ||
| T=1 | Quality | 7.430.2525 | 6.670.3139 | 7.570.2695 | 7.200.2599 | 7.200.2844 | 7.330.2722 |
| Diversity | 6.700.3371 | 6.230.3672 | 6.970.3415 | 6.500.3965 | 6.530.3358 | 6.630.3349 | |
| T=2 | Quality | 1.030.0328 | 7.370.2185 | 7.500.2656 | 1.300.2317 | 1.370.2645 | 6.030.3785 |
| Diversity | 1.070.0455 | 6.970.3696 | 7.100.3601 | 1.330.2325 | 1.370.2645 | 7.800.3812 | |
| T=3 | Quality | 1.030.0328 | 1.230.0905 | 5.800.3664 | 1.000.0000 | 1.030.0328 | 1.270.0935 |
| Diversity | 1.070.0455 | 1.270.0935 | 5.900.4555 | 1.000.0000 | 1.070.0455 | 1.300.1261 | |
C.3 Additional LLM-As-A-Judge Evaluation for Creative Writing
In addition to the AlpacaEval Creative Writing evaluation, we conducted our own LLM-As-A-Judge experiment comparing min_p against top_p sampling across multiple dimensions of text quality for Creative Writing. We also used this opportunity to test the performance of min_p on constrained/structured generation tasks. Our results provide strong evidence supporting min_pβs effectiveness, particularly at maintaining text quality across different temperature settings.
Specifically, we conducted a comprehensive evaluation using two language models of different scales:
β’
Llama-3.2-1B-Instruct (1B parameters)
β’
Mistral-7B-v0.1 (7B parameters)
C.3.1 Structured Generation Framework
To ensure consistent and comparable outputs, we implemented a structured generation approach using Pydantic schemas for the two models. We keep it simple as a baseline:
class CreativeStory(BaseModel):
themes: List[str]
writing_complexity: int = Field(ge=1, le=10)
short_story_text: str
The modelsβ outputs were constrained using lmformatenforcerβs JsonSchemaParser and transformers prefix token filtering, ensuring all generated stories followed the same structured format.
C.3.2 Creative Writing Task
We used three distinct creative writing prompts to evaluate generation quality:
1.
βWrite a story about a mysterious door that appears in an unexpected placeβ
2.
βWrite a story about an alien civilizationβs first contact with Earth from their perspectiveβ
3.
βWrite a story about a world where time suddenly starts moving backwardsβ
C.3.3 Sampling Parameters
We tested a comprehensive matrix of sampling parameters:
β’
Temperatures: [0.5, 1.0, 2.0, 3.0, 5.0]
β’
min_p values: [0.05, 0.1, 0.2]
β’
top_p values: [0.9, 0.95, 0.99]
For each combination, we generated stories using both min_p and top_p sampling methods, with all other parameters held constant.
C.3.4 Evaluation Methodology
Blind Comparison Setup
β’
For each comparison, stories from both sampling methods were randomly ordered as Response 1 or Response 2 (to mitigate position bias)
β’
The evaluation system was blind to which sampling method produced each response
β’
A GPT-4o model served as the judge, using a structured evaluation schema:
class LLMasJudge(BaseModel):
response1_creativity_score: Literal["0" to "10"]
response1_originality_score: Literal["0" to "10"]
response1_narrative_flow_score: Literal["0" to "10"]
response1_emotional_impact_score: Literal["0" to "10"]
response1_imagery_score: Literal["0" to "10"]
response2_[same metrics as above]
detailed_feedback: str
overall_winner: Literal["1", "2"]
Judge Configuration
β’
Model: GPT-4
β’
Temperature: 1.0 (to ensure consistent but non-deterministic evaluation)
β’
Structured output enforcement using OpenAIβs beta chat completions parse endpoint
β’
System prompt: βYou are an expert judge evaluating AI-generated creative writingβ
Evaluation Metrics
Each story was evaluated on five dimensions:
1.
Creativity: Novelty and uniqueness of ideas
2.
Originality: Innovative approach to the prompt
3.
Narrative Flow: Coherence and story progression
4.
Emotional Impact: Ability to evoke feelings
5.
Imagery: Vividness of descriptions
C.3.5 Data Collection and Analysis
β’
Results were logged to Weights & Biases for tracking and analysis, and all results will be published on github
β’
Each evaluation included:
β
Full generated stories from both methods
β
Detailed scores across all metrics
β
Judgeβs qualitative feedback
β
Randomized position tracking
β
Complete parameter configuration
β
This comprehensive setup allowed us to analyze the performance of min_p vs top_p sampling across different model sizes, temperatures, and parameter values while maintaining experimental rigor through structured generation and blind evaluation.
C.3.6 Results of constrained LLM-as-judge evaluation
Overall Performance
Our results show that min_p consistently outperforms top_p across all quality metrics:
| Metric | min_p | top_p | Difference |
| Creativity | 3.55 | 3.09 | +0.46 |
| Originality | 3.28 | 2.85 | +0.43 |
| Narrative Flow | 2.96 | 2.26 | +0.70 |
| Emotional Impact | 2.62 | 2.10 | +0.52 |
| Imagery | 2.98 | 2.36 | +0.62 |
Temperature Stability Analysis
A particularly notable finding is min_pβs superior performance at maintaining quality across different temperature settings. Table 17 compares performance across temperature settings from 0.5 to 5.0.
| Model | Metric | Low Temp (0.5) | High Temp (2.0) | Very High Temp (3.0) | |||
| min_p | top_p | min_p | top_p | min_p | top_p | ||
| Llama-1B | Creativity | 6.33 | 4.93 | 3.78 | 2.44 | 2.12 | 1.92 |
| Originality | 5.70 | 4.48 | 3.63 | 2.59 | 1.96 | 1.73 | |
| Narrative Flow | β | β | β | β | 1.19 | 1.04 | |
| Emotional Impact | β | β | β | β | 1.12 | 0.88 | |
| Imagery | β | β | β | β | 1.27 | 1.15 | |
| Mistral-7B | Creativity | 5.56 | 5.56 | 3.44 | 2.70 | 1.78 | 1.59 |
| Originality | 5.07 | 4.89 | 3.04 | 2.44 | 1.81 | 1.48 | |
| Narrative Flow | β | β | β | β | 0.96 | 0.74 | |
| Emotional Impact | β | β | β | β | 1.00 | 0.78 | |
| Imagery | β | β | β | β | 1.11 | 0.81 | |
Performance by Configuration (Temperature 1.0)
Table 18 shows detailed comparisons of different configurations at temperature 1.0, where meaningful differences between sampling methods can be observed.
| Model | Config | min_p | top_p | ||||||||
| Creat | Orig | Flow | Emot | Imag | Creat | Orig | Flow | Emot | Imag | ||
| Llama-1B | p=0.05/0.9 | 4.78 | 4.22 | 4.56 | 3.44 | 4.11 | 4.33 | 3.89 | 3.44 | 3.11 | 3.33 |
| p=0.1/0.95 | 5.00 | 4.67 | 4.11 | 3.89 | 3.67 | 5.44 | 4.78 | 3.33 | 3.00 | 3.11 | |
| p=0.2/0.99 | 5.33 | 5.33 | 4.56 | 3.78 | 5.00 | 4.44 | 4.11 | 3.56 | 4.11 | 4.33 | |
| Mistral-7B | p=0.05/0.9 | 5.44 | 5.11 | 5.22 | 4.67 | 4.67 | 4.67 | 4.44 | 4.67 | 4.11 | 5.11 |
| p=0.1/0.95 | 6.89 | 6.22 | 6.67 | 5.89 | 7.00 | 5.33 | 4.78 | 4.22 | 3.89 | 4.11 | |
| p=0.2/0.99 | 5.11 | 4.44 | 4.44 | 3.44 | 4.22 | 4.00 | 3.89 | 3.44 | 3.33 | 3.22 | |
The results demonstrate that min_p consistently outperforms top_p across various metrics, with especially notable improvements in narrative flow (+0.70) and imagery (+0.62). This advantage is maintained across temperature settings, with min_p showing particular strength at higher temperatures where coherence typically degrades. The optimal configuration appears to be min_p=0.1 with the Mistral-7B model, which shows substantial gains across all quality dimensions.
Appendix D Applications and Future Work
D.1 Real-World Applications of Min-p Sampling
Min- proves useful in domains where high-temperature incoherence was previously a bottleneck:
β’
Creative Writing: Enhances narrative generation by allowing higher temperatures without losing coherence, unlocking new capabilities for storytelling and poetry.
β’
Diverse Reasoning Paths: Facilitates problem-solving and brainstorming by generating varied outputs/reasoning paths via adaptive temperature. We note that such approaches are currently limited at higher temperature ranges which min- enables Dhuliawala etΒ al. (2024); xjdr & doomslide (2024); Zhang etΒ al. (2024). Wang Wang & Zhou (2024) found that even while solving basic arithmetic in GSM8K chain-of-thought (COT), diverse COT reasoning tokens outperform pure greedy COT decoding. This, alongside our results showing that min- sampling with higher temperature can outperform greedy decoding, suggests that Pareto optimal diversity -accuracy can surpass traditional deterministic approaches.
β’
Agent Training and Exploration: Recent research in reinforcement learning has begun utilizing min- sampling for generating high-quality and diverse training data for curious agents (Tajwar etΒ al., 2025). Specifically, they implemented min- sampling with temperature 1.5 and min- parameter 0.3 in their Llama-3.1-8B-Instruct model, finding that this configuration consistently produced more diverse trajectory data while maintaining higher real-time accuracy compared to alternative sampling methods. This approach enabled their agent to explore more effectively while maintaining solution coherence, demonstrating min-βs potential for improving reinforcement learning for agent exploration.
β’
Red-Teaming: Generates diverse samples for identifying vulnerabilities(Anurin etΒ al., 2024)
β’
Advanced Reasoning Models: Min- has shown particular promise with state-of-the-art reasoning models, with major providers like Unsloth (2025) explicitly recommending min- settings for their DeepSeek-R1 implementation. Their documentation specifies that "βmin- 0.05" helps "counteract very rare token predictions," particularly beneficial for their quantized 1.58-bit model, demonstrating how min- effectively balances token selection in both high-temperature creative settings and reasoning-intensive applications.
D.2 Limitations and Future Research Directions
While min- shows considerable promise, there are limitations and opportunities for future work:
Standard Deviation Sampling:
As of February 2025, we are aware of the Top-N (Tang etΒ al., 2024) method, which builds on min- by filtering tokens based on the standard deviation of the logit distribution. It defines a truncation threshold:
where and is the standard deviation of the logits.
We are actively exploring Min- sampling, a variation that replaces the mean-based threshold with a median-centered approach for robustness against skewed distributions. Min- derives its name from its use of the Z-score to normalize token logits before truncation, ensuring a more adaptive and information-theoretic selection criterion:
Here, each logit is transformed into a Z-score relative to the median rather than the mean, making the method more robust to heavy-tailed distributions and extreme outliers.
Both methods dynamically truncate low-probability tokens, but min- is designed to be more information-efficient in heavy-tailed logit distributions. More importantly, Min- extends beyond token probability truncation (which operates on a 0β1 scale); it can be applied to any ranked selection process, including discrete cases (such as expert routing in Mixture-of-Experts models or latent selection in Sparse Autoencoder training for mechanistic interpretability) and continuous cases (such as activation magnitude filtering or importance weighting in adaptive mechanisms). This provides a dynamic alternative to fixed Top- thresholds commonly used in Machine Learning.
Generalization to Other Models:
Our experiments focused on the Mistral and Llama 3 series models. Future work should explore the effectiveness of min- sampling with larger models and different architectures to assess its generalizability. Unsloth (2025) currently recommends using min- with DeepSeek-R1 by, the largest and most advanced open reasoning model available, suggesting promising applications across diverse model architectures beyond the ones weβve tested.
Hyperparameter Sensitivity:
The base probability threshold is a critical hyperparameter. Investigating methods for dynamically adjusting based on context or developing guidelines for optimal settings across various tasks could enhance performance. This was particularly challenging, as MAUVE hyperparameter sweeps are measured without temperature scaling on GPT2-XL (Pillutla etΒ al., 2023), and require pre-selection on exact hyperparameters. min-, however, is a novel method often used in conjunction with temperature scaling without extensive literature/experimental data. We discuss this further in SectionΒ A.2.
Combined Sampling Approaches:
While our own tests do not find benefits to combining truncation sampling methods, and there is little existing research into the matter, we acknowledge that combining truncation sampling approaches could be viable. For example, in theory implementing a min- threshold with a top-k threshold and choosing whichever is higher/lower could be dynamic. However, this is hard to test as we would have to isolate relative contributions of both methods, tune hyperparameters separately and account for benchmark variance among other things.
Theoretical Analysis:
A deeper theoretical understanding of why min- sampling performs better, particularly at high temperatures, could provide insights into the behavior of language models and guide the development of even more effective sampling strategies.
Applicability to Other Domains:
Extending min- to other generative tasks, such as code generation or multimodal models, could reveal broader applicability and benefits across different domains.
Research into high-temperature regimes:
High-temperature regimes have been underexplored relative to low-temperature regimes. Min- sampling hopes to unlock exploration, experimentation, and applications in such areas. Recent work in reinforcement learning has begun adopting min- for trajectory data generation (Tajwar etΒ al., 2025), specifically citing its benefits for generating high-quality and diverse training data. In their Llama-3.1-8B-Instruct model implementation, they selected min- sampling with temperature 1.5 and min- parameter 0.3 after observing that this configuration consistently generated diverse training data while maintaining higher real-time accuracy compared to other sampling methods. This demonstrates that principles from min- have valuable applications beyond text generation in areas like agent training and decision-making systems, where balancing exploration and coherence is crucial.
Human Evaluation Scope:
Our involved participants selecting pre-generated outputs. We note that min-βs popularity within the open source community for creative writing is interactive in nature; hence, we hope for adoption on interactive platforms such as the Chatbot Arena (Chiang etΒ al., 2024).
Combining uncertainty and CoT decoding methods:
Wang & Zhou (2024) found a significant correlation between the confidence/certainly level of the final answer token choice and correct scores on GSM8K CoT, and that promoting diverse lower-probability token choices encouraged generating chains of thought that were beneficial for reasoning, resulting in higher scores overall.
This mirrors our hypothesis that choosing high-certainty tokens enables more accurate final answers, while diverse token choices benefit intermediate reasoning steps. For example, we note that GPQA scores on Mistral 7B increased from to , but only with Temperature Only and min-. With the recent release of OpenAIβs O1 models, which leverage CoT methods at inference for advanced reasoning capabilities, we note several novel decoding methods that combine uncertainty and CoT sampling approaches to improve model reasoning in a simple manner, with minimal added overhead and architectural changes (Wang & Zhou, 2024; xjdr & doomslide, 2024). We aim to actively explore such methods in future work.